If the polynomial x3 + x2 – 2 is divided by x + 1, the remainder is 0. True O False
Answer #1

Plusz x +0²2 divided by 3+1 i.e. a = 1 P(-a)=P(-1) = (-1)3 +1-2=-141-2 =-2 Since P(-a) = P(-1270 Hence the given statement is